OpenStudy (anonymous):

what is the equation in standard form of a perpendicular line that passes through (5,1) the equation they gave me to answer the question is y-3=3(x+1)

OpenStudy (ranga):

First find the slope of the line: y-3=3(x+1) Can you find the slope?

OpenStudy (anonymous):

the slope is 3

OpenStudy (ranga):

Yes. The slope of the perpendicular line will be the negative reciprocal of this slope.

OpenStudy (anonymous):

what is the negative reciprocal of the slope

OpenStudy (ranga):

Slope of the perpendicular line = -1/3 Equation of a line in slope-intercept form is: y = mx + b m = -1/3 y = -1/3x + b It passes through (5,1). Put x = 5, y = 1 into the equation and solve for b. Then put the equation in the standard form.

OpenStudy (anonymous):

so is this how the answer's suppose to be 1=-1/3*5+b (* means multiply)

OpenStudy (ranga):

solve for b.

OpenStudy (ranga):

1 = -5/3 + b b = 1 + 5/3 = 8/3 y = -1/3x + 8/3 (this is the equation in slope-intercept form). They want the equation in standard form: multiply throughout by 3 3y = -x + 8 x + 3y = 8 is the equation of the perpendicular line in standard form.

OpenStudy (anonymous):

thank you : D

OpenStudy (anonymous):

what is the intercept of the perpendicular line this is right after the first question Ist asked

OpenStudy (ranga):

The y-intercept is the value for "b" we found earlier: 8/3 In y = mx + b, m is the slope and b is the y-intercept.

OpenStudy (anonymous):

what the x-intercept of the perpendicular line